The area of a square is numerically 77 more than the perimeter. Find the length of the side.
Answer by checkley79(3341)

s^2=4s+77
s^2-4s-77=0
(s-11)(s+7)=0
s-11=0
s=11 ans for the length of the sides.
Proof:
11^2=4*11+77
121=44+77
121=121